OpenStudy (anonymous):

What is the center of a circle with this equation: (x+4)^2 + (y-8)^2 = 81?

OpenStudy (shamil98):

The center is (h,k) (x-h)^2 + (y-k)^2 = r^2 (-4,8) are your values because (x-(-4)^2 + (y-+8)^2 = 9^2 becomes (x+4)^2 + (y-8)^2 = 81

OpenStudy (anonymous):

Thank you VERY much!